OpenStudy (anonymous):

solve the system 2x-2y+z=-1 x+3y+4z=11 3x+0y+2z=5

hero (hero):

Eliminate the z variable so that you'll be left with a system of two equations in variables x and y.

hero (hero):

You can do this easily by pairing the 1st and 3rd equations 2x-2y+z=-1 3x+0y+2z=5 Then by pairing the 2nd and 3rd equations: x+3y+4z=11 3x+0y+2z=5 Now, working with the 1st and 3rd equations, Multiply the 1st equation by 2, then subtract: 4x - 4y + 2z = -2 3x +0y + 2z = 5 After subtraction you get: x - 4y = -7 Working with the 2nd and 3rd equations, Multiply the 3rd equation by 2 then subtract: x + 3y + 4z = 11 6x + 0y + 4z = 10 After subtraction you get: -5x + 3y = 1 You're left with a system of two equations: x - 4y = -7 -5x + 3y = 1 Which you might be able to solve.

hero (hero):

For the system of 2 equations, you can multiply the first equation by 5 to get: 5x - 20y = -35 -5x + 3y = 1 Combining both equations together you get -17y = -34 34 = 17y Dividing both sides by 17 to get 2 = y From there, you can do back substitution to find x and z.

hero (hero):

The easiest way to solve this, however, is with matrices.

OpenStudy (anonymous):

x - 4y = -7 means x = 4y - 7 -5x + 3y = 1 means x = (1 - 3y) / -5 Equal both resulting equations, 4y - 7 = 3/5y - 1/5 Rearrange, and y = 2. If y = 2, x = 4(2) - 7 = 1. Thefore, y = 2 x = 1 Now, z, since 3x + 2z = 5, 3(1)+2z = 5 2z = 2, thefore z = 1. x = 1, y = 2, and z = 1.

OpenStudy (anonymous):

(I continued from @Hero 's first response) And he is very right about using matrices to make this easier.